Solve the following system of linear equations. (If the system has no solution, enter NONE for each answer. If the system is dependent, enter INFINITELY MANY for each answer.)
y=x+2
y=-x-6
x = ?
y = ?
Pls show me the steps on how to get the answer. thanks

Answer by rapaljer(4671)

y=x+2
y=-x-6

x+2=-x-6

Add +x to each side:
2x+2=-6
2x=-8
x=-4

Substitute back into the first equation:
y=x+2
y=-4+2
y=-2

Check in the second equation:
y=-x-6
-2=-(-4)-6
-2=4-6
It checks!!

For additional help with this topic, please see my own website. To find my website, you can click on my tutor name "Rapaljer" anywhere in algebra.com. There you will find a link that takes you to my Homepage. On the Homepage, look for the link "Basic, Intermediate, and College Algebra: One Step at a Time." Choose, "Basic Algebra", and look in "Chapter 4" for Section 4.07 "Systems of Equations." Here you will find my own non-traditional explanation that was written especially for students who have trouble with math. My own students all told me that it was much easier to understand than the traditional textbooks. Now, my entire curriculum from Basic Algebra to College Algebra is all posted on the website FREE!!

Dr. R^2, Retired
Seminole State College of Florida